# Language-specific Calibration for Pruning Multilingual Language Models

### v1

@@ 3.5 Implementation Details
4 Results
4.1 Perplexity
Which calibration language would lead to lower perplexity?
Table 1 presents the perplexity scores in different languages for pruned models. The column headers denote the evaluation languages, while the row headers indicate the calibration languages. Our findings reveal that there is no universal calibration language that consistently minimizes perplexity across all evaluation languages. In other words, the optimal calibration language to achieve lower perplexity varies depending on the target language being evaluated. Notably, as indicated by the darker diagonal pattern, calibrating on the target language itself consistently yields the lowest perplexity. The only exception occurs when evaluating Llama-3 pruned with Wanda on Chinese. In this case, the optimal calibration language is Russian, with Chinese as the second-best option. The perplexities of these two calibration languages are closely aligned, i.e., 24.4 for Chinese and 23.6 for Russian.
4.2 Downstream Task Performance
Which calibration language would lead to better downstream-task performance?
@@ 4.2 Downstream Task Performance
Table 2 shows the models’ performance on downstream tasks. The column headers denote the evaluation languages, while the row headers indicate the calibration languages. First, for both pruning methods, the choice of calibration language affects downstream task performance. This impact is particularly pronounced when pruning Llama-3 8B model with SparseGPT or Aya-23 models with either Wanda or SparseGPT, as evidenced by the greater color difference in the table, compared to pruning Llama-3 8B with Wanda (top quarter). For example, when evaluating Llama-3 8B model on Belebele in English, pruning with Wanda results in performance ranging from 58.5 with Spanish calibration to 65.0 with English calibration. In contrast, when pruning Llama-3 8B with SparseGPT the performance varies from 50.0 with Chinese calibration to 71.4 with English calibration.
@@
Unlike the perplexity evaluation in 1, calibration in the target language does not reliably result in good performance. For instance, on Belebele, the pruned Aya-23 8B model mainly achieves higher accuracy on evaluation languages other than on its calibration languages. Overall, of 50 evaluation groups22225 columns for Wanda and 25 for SpareGPT., in 26 group comparison cases, calibration in the target language yields the best performance, 8 for Wanda and 18 for SparseGPT (e.g. Arabic on ARC and MMLU). For the remaining cases, calibration in the target language achieves the second best or comparable performance. Therefore, calibrating using the target language often results in acceptable, though not consistently the best, performance. Given the standard deviations shown in Table 6 in Appendix B, calibration in the target language is a reasonable choice in terms of downstream performance.
@@
In terms of downstream language performance, the full-sized baseline models exhibit the strongest performance in English, followed by other Latin languages, with Russian next in line. Arabic and Chinese generally perform worse. Further, pruning can sometimes disrupt the original ranking of languages observed in the baseline models. For example, on the Belebele benchmark, the Llama-3 8B model achieves a baseline accuracy of for Arabic and for Chinese, but the pruned models reverse this trend and achieve an accuracy of for Arabic and for Chinese. That is, pruning can shift which languages the model performs best or worst on.
@@
Calibrating in an outlier language could enhance generalization for downstream in non-English?
@@
We include Swahili as an outlier language, which Llama-3 and Aya-23 do not include in the pre-training corpora. In column-wise comparison, the SW cell is consistently the darkest or relatively dark one. That is, in most cases, calibration in Swahili results in the worst or the second-worst performance across different calibration language strategies, given a downstream task.
Would calibrating in a similar language be helpful?
There was no consistent pattern based on the similarity of calibration-evaluation language pairs. For example, Latin language pairs such as English-Spanish (calibrating in English and evaluating in Spanish) or pairs from the same language family like English-German did not always perform well, while pairs with different writing systems, such as Russian-English or Spanish-Arabic, did not consistently perform poorly. On the other hand, calibration in a dissimilar language, i.e. Chinese, often results in particularly low accuracy across many tasks and evaluation languages, as demonstrated by the darker row of “ZH”.
To what extent, do the model and pruning method matter?
Between Llama-3 8B and Aya-23 8B, despite their identical decoder-only architecture, Table 2 highlights distinct performance patterns between the two models under investigation. With its larger vocabulary size in the six non-English languages, Aya-23 8B generally outperforms the Llama-3 8B model in most evaluation languages and tasks, both before and after pruning. Notably, Aya-23 8B experiences less performance drop after pruning but shows less stable results, often performing better on languages other than the one used for calibration. Overall, these results suggest that vocabulary size and pre-training data impact baseline performance and accuracy after pruning in a multilingual context.
Between Wanda and SparseGPT, Llama-3 8B’s performance degrades less after pruning with SparseGPT pruning, while the performance of the pruned Aya-23 8B model varies, that is, excelling with Wanda or SparseGPT depends on the task. Specifically, SparseGPT produces a distinct diagonal pattern for Aya-23 8B on ARC and HellaSwag, while Wanda often yields superior performance for calibration in non-target languages on Belebele and MMLU. No pruning technique consistently performs best in all tasks.
4.3 Impact of Model Sizes
To investigate the scaling impact towards pruning behavior, we repeat experiments in Table 6 on the larger Llama-3 70B and Aya-23 35B models, the results of which are reported in Appendix B. Overall, pruning performance increases with higher baseline accuracy of the full-sized models. However, we observe that the performance patterns and findings from the smaller models, mentioned above, do not consistently hold true on their bigger counterparts.
4.4 Open Domain Question Answering without Context
@@
AR-DE-EN-ES-RU | 30.4 | 34.9 | 69.5 | 40.2 | 35.7 | 33.2 | 30.5 | 59.0 | 69.1 | 75.4 | 63.0 | 64.9 | 30.4 | 65.8 | 37.3 | 42.6 | 46.9 | 44.6 | 41.2 | 40.4 | 38.1 | 42.8 | 51.8 | 45.8 | 40.6 | ||
AR-DE-EN-ES-RU-SW-ZH | 30.4 | 34.7 | 69.2 | 40.0 | 35.4 | 33.0 | 30.9 | 57.9 | 68.6 | 74.7 | 64.2 | 65.9 | 31.3 | 67.4 | 37.4 | 42.5 | 46.6 | 44.4 | 41.2 | 40.7 | 38.1 | 42.8 | 51.7 | 45.9 | 40.5 |
How does pruning impact the knowledge stored in multilingual LLMs?
Table 3 shows the open-domain question-answering performance of Llama-3 8B and Aya-23 8B across various calibration-evaluation language pairs. This “closed-book” task, MKQA, provides no context in the prompt, and relies on the model’s internal knowledge. To ensure a fair comparison between languages, the MKQA dataset is fully parallel and primarily features entity-based and structured “atomic” answer types.333There are a total of 6,758 evaluation samples after eliminating those with unanswerable and long, imprecise answers. First, we observe significant performance differences among the evaluation languages for the full-size models. While Latin languages perform best, Arabic and Chinese perform worst. Further, we notice a significant accuracy drop by pruning, even for performance in English. In summary, pruning substantially impacts the storage and retrieval of knowledge in a multilingual model across different languages.
4.5 Multiple Calibration Languages
Will incorporating more languages in the calibration, i.e.bilingual and multilingual calibration set, benefit the performance of pruned multilingual models?
We repeat the experiments but include more than one language in the calibration set. As English is the dominant language in pre-training data, we first test the combination of English and the target language for a bilingual calibration setup. We further experiment with including all seven languages in the calibration, i.e. multilingual setup. For all setups, the total calibration sample number remains the same, i.e. 128.
Comparing Table 4 with Table 2, we observe that bilingual and multilingual calibration sets generally outperform monolingual calibration for non-English target languages. For example, for Chinese on ARC, the seven-lingual calibration set, AR-DE-EN-ES-RU-SW-ZH, often yields higher accuracy than calibration in Chinese, i.e. the target language, across models and pruning methods. That is, AR-DE-EN-ES-RU-SW-ZH achieves 28.3 (Wanda, Llama-3 8B), 30.0 (SparseGPT, Llama-3 8B), 32.0 (Wanda, Aya-23 8B) and 30.9 (SparseGPT, Aya-23 8B), consistently higher or comparable to monolingual calibration set, i.e. 27.0, 30.1, 30.8 and 31.7 respectively. On the other hand, multilingual calibration pruning often degrades the downstream performance in English, particularly on ARC when pruning Aya-23 with SparseGPT and on MMLU when pruning Llama-3 with Wanda. Interestingly, the best bilingual combination does not necessarily contain the target language but English. For example, for Russian on ARC when pruning with Wanda, calibration in EN-ES and AR-EN leads to the optimal performance for Llama-3 8B and Aya-23 8B.
@@ 5 Analysis
Inspired by Xie et al. (2022), we explore how pruning affects language-specific and langauge-agnostic features in multilingual LLMs. Utilizing Low-rank Subspace for language-Agnostic Representations (LSAR) by Xie et al. (2022), we disentangle language-specific syntactic features from semantic language-agnostic features to investigate pruning behavior across different calibration languages for the final layer’s output embeddings. The dominance of language-specific elements within embeddings allows us to distinguish embeddings of various languages in low dimensional space through an SVD. Projecting back from this reduced space, using the most significant singular values, into the original embedding space provides the language-specific features . By subtracting from the original embedding , we isolate the language-agnostic features, .
Figure 1 compares the magnitude of the difference of the language-agnostic features (a) and language-specific features (b) of the final layer’s embeddings before and after pruning with SparseGPT on Llama-3 8B. First, the language-agnostic features remain similar among pruned models of different calibration languages when evaluated in different downstream languages (x-axis), whereas language-specific features changed noticeably as different languages were used for calibration. Particularly, as shown in sub-figure (b), calibrating in Germany results in the smallest change of language-specific features when evaluating in German, compared to calibration in other languages. We also observe similar patterns in Spanish and Swahili. We hypothesize that calibration in the target language helps retain language-specific features, which might benefit coherent and fluent text generation. This might be the reason behind the diagonal pattern observed in Table 1 in Section 1 that calibration in the target language results in the lowest perplexity. On the other hand, while these language-specific features enhance fluency, they are unlikely to improve the model’s understanding and reasoning capabilities, which are crucial for processing downstream tasks.
To support this assumption, we plot the layer-wise magnitude of mean differences between the hidden-states of the full-sized and pruned models in Figure 2 of Appendix B. Calibration in the evaluation language only benefits the first and few last layers that are directly responsible for processing and generating language-specific features, but offers minimal to no advantage for all other layers, including those associated with semantic understanding. This observation may elucidate why, as shown in Table 2, calibration in the target language does not lead to optimal performance on downstream tasks.
6 Recommendations
We summarize our findings into three recommendations for pruning multilingual language models to achieve optimal performance in specific languages.
Choice of Calibration Language For a given target language, it is advisable to calibrate the pruning of the multilingual model specifically in that target language, even if the primary pre-training language of the model is English. However, such calibration in the target language may result only in optimal perplexity and reasonable downstream performance, particularly in tasks requiring deep language understanding and reasoning, not just superficial coherence and fluency.
Choice of Pruning Method Between Wanda and SparseGPT, there is no better one across model families, tasks, and languages. In general, Wanda is recommended for Aya-23 models and SparseGPT for the Llama-3 models.
Evaluation The perplexity score that is often used for evaluating pruning performance can be misleading for downstream task performance. For that reason, we recommend directly testing on the downstream tasks instead. Moreover, due to the known phenomenon of potentially significant performance variance depending on the calibration set choice for certain tasks (Williams and Aletras, 2023), also demonstrated in the accuracy standard deviations in Table 6, we further suggest validating performance over multiple pruning runs and evaluation tasks.

### v3

@@ 4.2 Downstream Task Performance
How to select the calibration language to optimize performance in downstream tasks?
Table 2 shows the models’ performance on downstream tasks. The column headers denote the evaluation languages, while the row headers indicate the calibration languages. Given a downstream task in a specific language, we analyze the results column-wise; the lighter the color in a column, the better the performance on the specified downstream task.
@@
First, for both pruning methods, the calibration language affects downstream task performance. This impact is particularly pronounced when pruning Llama-3 8B model with SparseGPT or Aya-23 models with either Wanda or SparseGPT, as evidenced by the greater color difference in the table, compared to pruning Llama-3 8B with Wanda. For example, when evaluating Llama-3 8B model on Belebele in English, pruning with Wanda results in performance ranging from 58.5 with Spanish calibration to 65.0 with English calibration. In contrast, when pruning Llama-3 8B with SparseGPT the performance varies from 50.0 with Chinese calibration to 71.4 with English calibration.
@@
Unlike the perplexity evaluation in Table 1, calibration with the target language does not reliably result in good performance. For instance, on MMLU, the pruned Llama-3 8B model mainly achieves higher accuracy on evaluation languages other than on its calibration languages. Overall, there are 50 evaluation groups, 25 columns for Wanda and 25 for SpareGPT, in 26 comparison cases. Calibration with the target language yields the best performance, 8 for Wanda and 18 for SparseGPT (e.g. Arabic on ARC and MMLU). For the remaining cases, calibration with the target language achieves the second-best or comparable performance. Therefore, calibrating using the target language typically results in acceptable, though not consistently the best performance. Considering the standard deviations shown in Table 7 in Appendix E, calibration with the target language is a reasonable choice out of downstream performance.
@@
In terms of downstream performance across different languages, the full-sized baseline models exhibit the strongest performance in English, followed by other Latin languages, with Russian next in line. Arabic and Chinese downstream tasks generally are the most challenging for models. Further, pruning can sometimes alter the original ranking of languages observed in the baseline models. For example, on the Belebele benchmark, the Llama-3 8B model achieves a baseline accuracy of for Arabic and for Chinese, but the pruned models reverse this trend and achieve an accuracy of for Arabic and for Chinese. That is, pruning can shift which languages the model performs best or worst on.
@@
Calibrating on an outlier language or a similar one could benefit downstream tasks in non-English?
@@
We include Swahili as an outlier language, which is out-of-domain for the model as it is not included in the pre-training corpora of Llama-3 and Aya-23. In column-wise comparison, the SW cell is consistently the darkest or relatively dark one. That is, in most cases, calibration with Swahili results in the worst or the second-worst performance across different calibration language strategies, given a downstream task.
@@
Furthermore, there is no consistent pattern related to the similarity of calibration-evaluation language pairs. For instance, Latin language pairs such as English-Spanish (calibrating in English and evaluating in Spanish) or pairs from the same language family, like English-German, do not always yield optimal performance. Conversely, pairs with different writing systems, such as Russian-English or Spanish-Arabic, do not consistently perform poorly. On the other hand, calibration with a dissimilar language, i.e. Chinese, often results in particularly low accuracy across many tasks and evaluation languages, as demonstrated by the darker row of “ZH”. In summary, we observe no benefit from calibrating with an outlier language or a similar language.
Does the model or the pruning method matter?
Between Llama-3 8B and Aya-23 8B, despite their similar decoder-only architecture, Table 2 highlights distinct performance patterns between the two models under investigation. With its larger vocabulary size in the six non-English languages, Aya-23 8B generally outperforms the Llama-3 8B model in most evaluation languages and tasks, both before and after pruning. Notably, Aya-23 8B experiences less performance drop after pruning but shows less stable results, often performing better on languages other than the one used for calibration. Overall, these results suggest that vocabulary size and pre-training data impact baseline performance and accuracy after pruning in a multilingual context.
Between Wanda and SparseGPT, Llama-3 8B’s performance degrades less after pruning with SparseGPT pruning, while the performance of the pruned Aya-23 8B model varies, that is, excelling with Wanda or SparseGPT depends on the task. Specifically, SparseGPT produces a distinct diagonal pattern for Aya-23 8B on ARC and HellaSwag, while Wanda often yields superior performance for calibration with non-target languages on Belebele and MMLU. No pruning technique consistently performs best in all tasks.
4.3 Open Domain Question Answering without Context
How does pruning impact the knowledge stored in multilingual LLMs?
Table 3 shows the open-domain question-answering performance of Llama-3 8B and Aya-23 8B across various calibration-evaluation language pairs. This “closed-book” task, MKQA, provides no context in the prompt, and relies on the model’s internal knowledge to generate answers. To ensure fair cross-languages comparisons, the MKQA dataset is fully parallel and primarily consists of entity-based and structured “atomic” answer types. See Appendix C for further details. First, we observe significant performance differences among the evaluation languages for the full-size models. While Latin languages perform best, Arabic and Chinese perform worst. Further, we notice a significant accuracy drop by pruning, even for performance in English. In summary, pruning substantially impacts the storage and retrieval of knowledge in a multilingual model across different languages.
4.4 Multiple Calibration Languages
@@
Five languages | 30.4 | 34.9 | 69.5 | 40.2 | 35.7 | 33.2 | 30.5 | 59.0 | 69.1 | 75.4 | 63.0 | 64.9 | 30.4 | 65.8 | 37.3 | 42.6 | 46.9 | 44.6 | 41.2 | 40.4 | 38.1 | 42.8 | 51.8 | 45.8 | 40.6 | ||
Seven languages | 30.4 | 34.7 | 69.2 | 40.0 | 35.4 | 33.0 | 30.9 | 57.9 | 68.6 | 74.7 | 64.2 | 65.9 | 31.3 | 67.4 | 37.4 | 42.5 | 46.6 | 44.4 | 41.2 | 40.7 | 38.1 | 42.8 | 51.7 | 45.9 | 40.5 |
Using more languages in the calibration, i.e. bilingual and multilingual calibration set, will benefit the performance of pruned multilingual models?
We repeat the experiments but include more languages in the calibration set. As English is the dominant language in pre-training data, we first test the combination of English and the target language for a bilingual calibration setup. We further experiment with including all seven languages in the calibration, i.e. multilingual setup. For all setups, the total calibration sample number remains the same, i.e. 128.
Comparing Table 8 with Table 2, we observe that bilingual and multilingual calibration sets generally outperform monolingual calibration for non-English target languages. For example, for Chinese on ARC, the seven-lingual calibration set, AR-DE-EN-ES-RU-SW-ZH, often yields higher accuracy than calibration with Chinese, i.e. the target language, across models and pruning methods. That is, AR-DE-EN-ES-RU-SW-ZH achieves 28.3 (Wanda, Llama-3 8B), 30.0 (SparseGPT, Llama-3 8B), 32.0 (Wanda, Aya-23 8B) and 30.9 (SparseGPT, Aya-23 8B), consistently higher or comparable to monolingual calibration set, i.e. 27.0, 30.1, 30.8 and 31.7 respectively. On the other hand, multilingual calibration pruning often degrades the downstream performance in English, particularly on ARC when pruning Aya-23 with SparseGPT and on MMLU when pruning Llama-3 with Wanda. Interestingly, the best bilingual combination does not necessarily contain the target language but English. For example, for Russian on ARC when pruning with Wanda, calibration with EN-ES and AR-EN leads to the optimal performance for Llama-3 8B and Aya-23 8B.
@@
4.5 Impact of Model Sizes
To investigate the scaling impact towards pruning behavior, we repeat experiments in Table 7 on the larger Llama-3 70B and Aya-23 35B models, the results of which are reported in Appendix E. Overall, pruning performance increases with higher baseline accuracy of the full-sized models. However, we observe that the performance patterns and findings from the smaller models, mentioned above, do not consistently hold true on their bigger counterparts.
4.6 Quantization
We further explore the impact of the calibration language in quantization on downstream task performance across different languages. We use GPTQ (Frantar et al., 2023) to quantize weights to 4 bits with a group size of 128 and 8 bits with a group size of 128 (equivalent to 50% sparsity in pruning) on LLaMA-3-8B. We follow our pruning setup for downstream tasks and languages for calibration and evaluation. The results are present in Table 14 and 15 in Appendix F, revealing several key findings that are consistent with our findings on pruning: (1) calibrating with the target language tends to yield better downstream performance; (2) pruning can alter which languages the model performs best or worst on; and (3) calibrating with an outlier or a linguistically similar language does not provide any notable advantage.
@@ 5 Analysis
In reviewing Table 1 and Table 7 together, we poist that calibrating on the target language effectively preserves language-related features but not language-agnostic features, such as knowledge memories and reasoning abilities. This implies that the smallest weights or weights of smallest activations contribute to the model’s nuanced reasoning or knowledge memorization and retrieval processes. The rationale here is that both SNR and pruning error measure the extent of model change after pruning, which is dominated by large values, that pruning aims to preserve. On the other hand, perplexity reflects the overall language modeling capability. The similar diagonal patterns across the three metrics, SNR, pruning error and perplexity, as shown in Table 1, suggest that calibrating on the target language, i.e. identifying and pruning the smallest values, preserves the next-token prediction capability (perplexity) in the target language, along with the optimal pruning performance in terms of pruning the smallest values (SNR and pruning errors).
However, optimal performance in downstream tasks requires more than just surface-level language modeling; it also depends on robust reasoning capabilities. The suboptimal performance observed in downstream tasks may be attributed to the degradation of language-agnostic reasoning capabilities. This phenomenon suggests that removing the weights or weights of the smallest activations might inadvertently compromise the model’s ability to process reasoning, underscoring the reasoning role of these pruned weights, i.e. the smallest weights or weights of the smallest activations.
To validate our hypothesis, we investigate the internal changes of pruned models on three different levels: subspace level, matrix level, neuron level (columns in a matrix, followed by non-linearity). Previous studies on multilingualism of LLMs have sought to separate language-specific features from language-agnostic features, either at the neuron level (Tang et al., 2024; Zhao et al., 2024; Wang et al., 2024) or from an extracted subspace perspective (Xie et al., 2022). To comprehend the underlying mechanisms driving the disparities in pruning metrics and downstream performance when using different calibration languages, we examine the shifts in language-specific and agnostic features and neurons induced by the pruning process.
5.1 language-specific Subspace Representations
The core idea of Low-rank Subspace for language-Agnostic Representations (LSAR) by Xie et al. (2022) is to construct a mean embedding matrix by concatenating L language embeddings. Subsequently, LSAR decomposes into a vector representing shared signals across languages and a matrix specifying a low-rank subspace on which different languages express different linguistic signals. This decomposition process is achieved via singular value decomposition on solving: with the orthogonality constraint. represents the coordinates of language-specific signals along the subspace’s components, and is a vector of all ones. We can extract language-specific features from a sentence embedding using by projecting into and back from the low-rank, language signal retaining subspace, . See Appendix D.2 for more details and the experiment setup.
We employ LSAR to decomposite the token-wise averaged embedding output of each layer of SparseGPT-pruned Llama-3 8B into language-specific and language-agnostic features, and estimate their changes after pruning. Figure 1 shows the layer-wise magnitude of differences ( magnitude) of (a) language-agnostic features and (b) language-specific features after pruning. From a high-level perspective, the greater magnitude suggests the greater changes in terms of hidden state after pruning, thus the greater pruning error and the worse pruning performance. Unlike the pruning error presented in Table 1, here we separately analyze the pruning error for language-agnostic and language-specific features. See Figure 5 in Appendix E for layer-wise full plots.
First, calibrating on the target language leads to relatively smaller pruning errors in terms of language-specific features as shown in the sub Figure (b) in Figure 1. This phenomenon is observed across layers, particularly noticeable in layer 32, as pinpointed by stars in Figure 1. This potentially explains the findings in Section 4.1 that calibration on the target language leads to the lowest perplexity, which is associated with a robust language-specific linguistic modeling capability. On the other hand, as indicated by the relatively flat horizontal lines across languages and layers in sub Figure (a), the pruning error on the language-agnostic features remains similar regardless of the calibration languages. This pattern helps explain the sub-optimal downstream task performance observed in Table 7, where no single calibration language consistently yields optimal performance across downstream tasks, including cases where the calibration is performed on the target language. That is, the selection of calibration language is unlikely to impact the language-agnostic features associated with understanding, reasoning, and knowledge retrieval. In contrast, the preservation of language-specific features, such as modeling capabilities related to perplexity, depends on the selection of calibration languages.
Second, the middle layers (as shown in the second and third columns in Figure 1) exhibit greater magnitude on language-agnostic feature representations and smaller magnitude on language-specific feature representations. This indicates that pruning errors can be predominantly attributed to the pruning errors on language-agnostic features, with less pruning error arising from language-specific features. Therefore, we conclude pruning generally affects language-agnostic features—potentially associated with reasoning and knowledge storage—more significantly than it impacts language-specific features.
5.2 Pruning Mask Similarity
To better understand the internal changes within the model after pruning, we conduct a matrix-level analysis by calculating the Intersection over Union (IoU) of pruning masks across different calibration sets, obtaining a measure of pruning mask similarity. The formal expression and details explaning IoU are presented in Appendix D.3. We use the pruning masks from the Llama-3 8B model pruned with SparseGPT for 50% unstructured sparsity. To reduce calibration set-dependent noise as prevalent in the downstream tasks, we first compute the intersection of pruned neuron indices across three pruned models calibrated with different seeds but in the same language . This intersection represents more stable neuron indices.
The IoU of the left side of Figure 2 depicts the proportion of with respect to all pruned neuron indices. It reveals high pruning mask similarity in the attention query, key and value in the first layer, while the attention output projection varies more significantly, especially after the 20th layer. This suggests that pruning struggles to consistently identify essential neurons in the attention output projection, partly responsible for the models reasoning capabilities.
The right plot compares the IoU of intersected neuron indices from the left plot across calibration in German, English, and Chinese. Notably, the attention query, key, and value in the first layer consistently achieve high IoUs above 0.95 across all languages, indicating that these components handle inputs similarly, irrespective of language differences. However, the attention output and FFN down projection show lower IoU, especially in early layers, with similarity peaking in middle layers (3rd to 15th) before decreasing again in later layers.
This pattern suggests that the attention output and FFN down projection in early and late layers handle language-specific signals, while middle layers process language-agnostic signals, supporting Figure 1. In other words, early layers focus on language comprehension, middle layers on language-independent reasoning, and later layers on generating language-specific predictions. This aligns with Zhao et al. (2024), who proposes that LLMs first comprehend queries by converting multilingual inputs into English in the early layers, reason in English in the intermediate layers, and then generate responses aligned with the input language in the final layers.
5.3 Language-specific Neurons
This section investigates neuron-level changes after pruning using Language Activation Probability Entropy (LAPE) as introduced by Tang et al. (2024). We focus on neurons of the up projections of the feed-forward layers (FFNs) followed by the non-linearity. A neuron is considered activated when the non-linearity output is greater than zero. LAPE measures the likelihood of individual neurons activating across different language inputs, taking neurons with high activation probability for one or two languages and lower probabilities for all others (i.e. low LAPE score) as language-specific. Details on the LAPE calculation and experiment settings are provided in Appendix D.4.
In Figure 3, each bar and boxplot corresponds to the same group of neurons, comprising 2% of the total neuron population. Neurons in each plot are sorted in ascending order based on their LAPE score in the full-sized model, enabling a comparison of changes in LAPE scores after pruning. We only plot the neuron groups of the 30% lowest LAPE score due to the rapidly diminishing variance for boxplots of higher LAPE scores. This suggests a lower impact of pruning on the activation frequency of high LAPE neurons, i.e. language-agnostic neurons. On the other hand, for the neuron group with the lowest LAPE scores (the leftmost boxplot), the higher the model-sparsity is, the longer whiskers (higher entropy variance) these neurons have. This indicates that pruning introduces LAPE noise, shifting the LAPE score distribution and creating new language-specific (low LAPE) and agnostic (high LAPE) neurons. The distribution shift of language-agnostic neurons in FFNs may contribute to performance degradation in downstream tasks. This hypothesis aligns with previous causal tracing studies, which have frequently identified FFNs are crucial for knowledge retention and retrieval (Meng et al., 2022; 2023).
Moreover, by analysis of boxplots and the bottom-right plot together, we find that the neuron group of the lowest LAPE has the lowest activation probabilities, while neurons with the highest LAPE scores activate more frequently. This indicates that pruning struggles to retain low activation probability of language-specific neurons. We also examine the absolute neuron output magnitude for these neglected neurons. The bottom plot in Figure 3 shows that despite the low activation probability, low LAPE neurons tend to have high output magnitudes. Pruned models show no significant differences in output magnitudes, leading us to conclude that pruning retains neuron output magnitudes in FFN modules but may fail to preserve activation frequency.